# The Physics of Price Discovery: Deconvolving Information, Volatility, and the Critical Breakdown of Signal during Retail Herding

### v1

@@ 4.3 The Critical Breakdown: Regime-Dependent Impact
Additionally, note that half-lives extend during herding for all investor types. This suggests impact persists longer when the market is in a critical state, consistent with reduced efficiency and slower price adjustment.
@@ 5.1 The Dual-Channel Market Structure
Our results reveal a fundamental duality in how order flow information is encoded. $S_{\text{MC}}$ serves as the “true signal” channel, capturing directional information about fundamental value. This is particularly evident for Institutional investors, where the 14-fold improvement after volatility adjustment demonstrates that $S_{\text{MC}}$ isolates pure information flow.
In contrast, $S_{\text{TV}}$ operates primarily as a “volatility proxy” channel. While useful for measuring execution costs and temporary price impact, it confounds information with liquidity conditions. A high $S_{\text{TV}}$ period may indicate informed trading, or it may simply reflect a low-liquidity regime where any trade moves prices.
5.1.1 Aggregation and Heterogeneity
The divergence between our Global (Table 1) and Conditional (Table 4) results highlights an important microstructure reality: price impact is highly heterogeneous across firm size. Table 1 and Panel B of Table 4 use pooled aggregation, which implicitly weights observations by data availability and market presence, effectively capturing the “market-cap weighted” physics. Here, Institutional investors show a positive time-weighted impact (+0.0025+0.0025), consistent with their role in efficient price discovery for large-cap stocks.
However, Panel A of Table 4 uses by-stock aggregation (equal weighting), reflecting the experience of the “typical” stock. The deeply negative impact during herding (0.0447-0.0447) suggests that for the average stock—likely smaller and less liquid—institutional order flow becomes toxic when retail investors herd. Crucially, both aggregation methods confirm the same physical breakdown: whether measured by the equal-weighted typical stock (Panel A) or the liquidity-weighted market aggregate (Panel B), institutional price impact significantly deteriorates during herding regimes, differing only in the degree of reversal. This duality suggests that while the market in aggregate remains somewhat efficient, the cross-section of stocks suffers a severe breakdown in price discovery during retail contagion events.
@@ 5.1.1 Aggregation and Heterogeneity
5.2 The Critical State of Retail Herding
The near-critical branching ratio (0.998) for retail trading surges is a remarkable finding. In the language of statistical physics, this places the market at a “critical point”—the boundary between ordered and disordered phases. Systems near criticality exhibit several characteristic features that we observe in our data:
Avalanches: A single event can trigger a cascade of subsequent events, leading to power-law distributed event sizes. This is consistent with documented “retail frenzies” in specific stocks.
Long-range correlations: Near-critical systems exhibit correlations over long timescales. The extended half-lives during herding periods (Table 4) support this.
Phase transitions: Small parameter changes can trigger qualitative shifts in system behavior. Here, the phase transition is in the information transmission mechanism itself. When retail intensity crosses the critical threshold, the market transitions from an efficient (signal-dominated) state to an inefficient (noise-dominated) state.
The regime flip in institutional impact (from +0.0047 to –0.0447) is the empirical signature of this phase transition. During normal periods, institutional flows are embedded in prices as they should be. During herding episodes, the same flows generate opposite effects—price initially moves in the direction of the flow, but subsequently reverses, yielding negative cumulative impact.
Why does this occur? The noise barrier hypothesis: when retail herding is intense, the market becomes saturated with noise trades. Institutional orders, even if fundamentally motivated, are “camouflaged” within the noise and cannot be distinguished by other market participants. As the herding episode resolves, prices revert, causing the initial institutional-driven move to reverse. From an institutional perspective, their information was “trapped” by the noise barrier.
@@ 5.3 Practical Implications
  • Develop real-time monitoring of branching ratios for retail trading activity.
  • High branching ratios (approaching 1) may warrant circuit breakers or cooling-off periods to prevent market breakdown.
@@ 5.4 Limitations and Future Research
Several limitations warrant discussion. First, our analysis focuses on a single market (Korea). While Korea offers excellent data on investor classification, generalization to other markets requires verification. Markets with different institutional structures (e.g., more algorithmic trading in the U.S.) may exhibit different herding dynamics.
Second, we analyze daily frequency data to avoid microstructure noise. However, intraday dynamics may differ. High-frequency herding episodes (e.g., within-day flash crashes) could exhibit even more extreme breakdowns in information transmission. Future work should extend the regime-conditional deconvolution framework to intraday horizons.
Third, our regime classification treats herding as endogenously determined by past trading activity (Hawkes model). However, some herding episodes may be triggered by exogenous news events (e.g., pandemic announcements). Distinguishing exogenous vs. endogenous herding triggers would enrich the analysis.
Fourth, while we document the regime flip empirically, the precise micro-foundations remain an open question. Agent-based models incorporating heterogeneous investors (informed institutional, herding retail) could provide theoretical grounding. Does the breakdown occur due to adverse selection (informed traders withdraw), inventory constraints (market makers reduce quotes), or attention constraints (investors cannot process information during noise saturation)?
Finally, the near-critical branching ratio (0.998) invites theoretical investigation. Is this an equilibrium outcome or a regulatory byproduct? Do markets naturally evolve toward criticality (as suggested by self-organized criticality literature (Bak et al.,, 1987)), or are they pushed there by regulatory constraints on trading?
@@ 6 Conclusion
This paper demonstrates that price discovery is not a mechanical process but a state-dependent phenomenon governed by the interplay of information and noise. We make three contributions to understanding this physics.
First, we resolve the normalization debate through information geometry. Market Cap Normalization ($S_{\text{MC}}$) isolates the pure directional signal of order flow, particularly for institutional investors where its informational content is 14 times greater than Volume Normalization ($S_{\text{TV}}$) after controlling for volatility. $S_{\text{TV}}$ primarily captures volatility regimes, not information.
Second, we establish the physical roles of market participants through impulse response analysis. Foreign and institutional investors act as architects of price discovery, exhibiting positive, persistent impact. Individual investors act as liquidity providers with negative cumulative impact, their buying pressure marking local price maxima.
Third, and most significantly, we document a critical breakdown in market efficiency during retail herding episodes. Using Hawkes process analysis, we show that individual trading surges exhibit near-critical self-excitation (branching ratio 0.998). During high herding periods, institutional price impact reverses from positive to negative—a 9.44-fold flip. This regime transition suggests the market enters an inefficient state where noise overwhelms signal, temporarily disabling information transmission.
These findings reframe market efficiency as a state variable, not a structural constant. When retail herding intensity is low, markets function efficiently: informed flows embed information, and prices discover value. When herding crosses the critical threshold, markets become informationally opaque: even sophisticated flows generate noise-like patterns, and price discovery breaks down.

### v2

@@ 4.2 The Critical Breakdown: Regime-Dependent Impact
Additionally, note that half-lives extend during herding for all investor types. This suggests impact persists longer when the market is in a near-critical state, consistent with reduced efficiency and slower price adjustment.
4.3 Information-Theoretic Validation: Irreversibility and Memory
We provide two independent validation tests that confirm and extend the physical segmentation documented above.
4.3.1 Entropy Production: Information vs. Impact
Table 5 reports the entropy production rate (EPR) for each investor type’s joint flow–return dynamics. The results differentiate the physical roles of market participants.
Table 5: Entropy Production Rate by Investor Type
Investor Type EPR 95% CI Perm. $p$-value Interpretation
Foreign 0.372∗∗∗ [0.218, 0.667] << 0.001 Irreversible (information)
Institutional 0.084 [0.077, 0.256] 0.405 Reversible (temporary impact)
Individual 0.142∗∗∗ [0.125, 0.519] << 0.001 Irreversible (cascade momentum)
Note: Entropy production rate computed on ternary-symbolized daily aggregate flow and return series (1,258 trading days, 9 joint states). Significance via permutation test (200 shuffles). ∗∗∗: $p<0.01$.
Foreign investors exhibit the highest EPR (0.372, $z=23.3$, $p<0.001$), indicating their flow–return dynamics are strongly time-irreversible. When foreign investors buy today and returns follow tomorrow, the temporal sequence looks fundamentally different when played backwards. This is the hallmark of permanent information injection: foreign flows embed directional content that persistently moves prices, consistent with their positive cumulative impulse response (Table 2).
Institutional investors show EPR indistinguishable from the null (0.084, $p=0.405$), indicating time-reversible dynamics. The two results are complementary: institutional flows move prices in the right direction on net (positive cumulative kernel in Table 2), yet the temporal pattern of flow and return is statistically symmetric—consistent with gradual portfolio rebalancing that nudges prices toward fundamentals without the sharp lead-lag signature of informed trading.
Individual investors exhibit moderate but significant EPR (0.142, $z=4.8$, $p<0.001$). Importantly, this irreversibility reflects a different mechanism than foreign information: given the overwhelming sell-side composition of surge events (Table 3), the irreversibility captures momentum from self-reinforcing feedback loops rather than fundamental information. The one-directional cascade creates flow–return dynamics that do not time-reverse—not because information is being injected, but because the panic itself generates irreversible momentum.
The EPR ordering (Foreign $\gg$ Individual >> Institutional) confirms the investor-type segmentation established in Section 4.1, with an important nuance: individual irreversibility reflects cascade momentum rather than information.
Robustness: The ordering is preserved under binary median symbolization (Foreign 0.074, Individual 0.072, Institutional 0.001; $p<0.001$ for Foreign and Individual). Quintile-5 symbolization (25 joint states) exceeds the resolution supported by 1,258 observations and loses discriminatory power across all investor types.
Figure 3: Entropy Production Rate by Investor Type. Bar heights show point EPR estimates with 95% block bootstrap confidence intervals. Foreign flows exhibit the highest time-irreversibility ($p<0.001$), consistent with permanent price impact. Institutional flows are statistically reversible ($p=0.405$). Individual flows show moderate irreversibility from flow-return feedback. Significance stars from permutation test: $^{***}p<0.01$.
4.3.2 Hawkes Memory: Systematic Programs vs. Sporadic Panic
To characterize the temporal structure of trading surges, we analyze inter-event clustering using the estimated Hawkes decay rate $\hat{\beta}$ as the natural timescale. Table 6 reports the results.
Table 6: Hawkes Memory Model: Clustering and Persistence by Investor Type
Investor $\beta^{-1}$ (days) Mean Cluster Max Cluster Frac. Isolated Gap CV Memory Depth
Foreign 11.9 4.7 44 0.33 1.80 4
Institutional 17.7 9.2 106 0.39 2.06 20
Individual Hawkes estimation failed (explosive dynamics)
[…]
Individual investors are the critical contrast: Hawkes parameter estimation fails entirely for their surge events. This reflects a structural property of individual surge dynamics: they are too explosive for the stationary Hawkes model. This is fully consistent with the unconstrained branching ratio of 1.09 documented in Section 4.2 and Appendix A.1—individual herding represents non-stationary, explosive feedback that cannot be characterized by a stable self-exciting process. Unlike foreign investors who execute systematic programs with predictable temporal structure, individual investors generate sporadic panic cascades whose timing is fundamentally unpredictable.
[…]
Together, the EPR and memory analyses provide independent confirmation that the investor-type segmentation (Section 4.1) extends to the level of temporal dynamics, with each investor type exhibiting qualitatively distinct patterns of irreversibility and clustering.
4.4 Mechanism: Noise Barrier vs Liquidity Withdrawal
Our size-stratified analysis reveals that institutional price impact deteriorates during herding episodes, particularly in small-cap stocks. But what is the physical mechanism driving this breakdown? We test two competing hypotheses using microstructure data.
The Noise Barrier hypothesis posits that during herding, retail trading volume explodes while market depth remains stable. The flood of uninformed order flow dilutes the institutional signal, making it harder to distinguish informed trades from noise. This predicts signal failure should correlate with high volume but stable bid-ask spreads.
The Liquidity Withdrawal hypothesis posits that during herding, market makers pull quotes in fear, widening spreads and reducing depth. This causes price dislocation as institutional orders face deteriorating execution conditions. This predicts signal failure should correlate with widening spreads and reduced depth.
We compute daily microstructure proxies for 126 identified high-herding days and 126 matched control days (252 days total). For each stock-day, we compute:
  • Realized volatility proxy: cross-sectional standard deviation of returns
  • Liquidity proxy: turnover (higher turnover indicates better liquidity)
  • Depth proxy: volume / realized volatility (Amihud-style illiquidity inverse)
We measure institutional signal efficacy [^1] Our liquidity proxy (turnover) is volume-based, which limits the sharpness of the Noise Barrier versus Liquidity Withdrawal distinction. Absent direct bid-ask spread data, our test identifies the relative explanatory contribution of volume-driven versus depth-driven proxies rather than cleanly separating the two mechanisms. as $S^{\text{inst}}_{\text{MC}}\times R_{t}$ (positive indicates aligned signal, negative indicates adverse selection). We then regress signal efficacy against these liquidity metrics, controlling for regime.
Table 7 presents the regression results.
Table 7: Mechanism Testing: Noise Barrier vs Liquidity Withdrawal
Model $R^{2}$ Key Coefficient Interpretation
Model 1: Noise Barrier 0.0093 Volume (–), Volatility (+) Volume dilutes signal
Model 2: Liquidity Withdrawal 0.0050 Liquidity (+), Depth (+) Weak liquidity effect
Model 3: Combined 0.0106 Both significant Volume contributes more
Model 4: Interaction 0.0040 Herding $\times$ Liquidity Regime-dependent
Note: N = 570,788 observations (294,759 herding, 276,029 normal). Standard errors are heteroskedasticity-robust (HC3). Both models explain <<1% of variance in absolute terms; in relative terms, Model 1 explains 86% more variance than Model 2.
Both models explain very little variance in absolute terms ($R^{2}<0.01$), so these results should be interpreted cautiously. Daily microstructure proxies—particularly without direct bid-ask spread data—capture only a small fraction of institutional signal efficacy. Within this limited explanatory power, the Noise Barrier model ($R^{2}=0.0093$) explains roughly twice as much variance as the Liquidity Withdrawal model ($R^{2}=0.0050$), offering suggestive but not conclusive evidence that volume explosion may be the more relevant channel. Sharper discrimination would require tick-level spread and depth data, which we leave to future work.
We also tested whether critical slowing down indicators (rolling autocorrelation and variance of retail flow) could predict regime transitions. Granger causality tests and ROC analysis both fail to detect predictive power (Appendix A.8), consistent with herding episodes emerging as sudden explosive shifts rather than gradual approaches to criticality.
@@ 5.1 The Dual-Channel Market Structure
The dual-channel segmentation established in Section 4.1 and independently validated in Section 4.3 has practical consequences that depend on the aggregation lens. Pooled estimates in Table 4 reflect the experience of liquid large-cap stocks, while equal-weighted estimates reflect the typical (smaller) firm—a divergence with direct implications for how practitioners should interpret aggregate efficiency metrics.
@@
5.2 Near-Critical Dynamics and Market Segmentation
The size-dependent patterns in Table 4 connect to a broader literature. Amihud, (2002) demonstrate that illiquidity effects concentrate in small-cap stocks, while Chordia et al., (2008) show that information transmission varies systematically with firm size. Our results extend these findings: near-critical herding dynamics selectively degrade efficiency where liquidity buffers are thinnest.
The amplification of positive institutional impact in large-cap stocks during herding may reflect increased arbitrage opportunities as sophisticated investors exploit temporary mispricings. Mid-cap stocks occupy an intermediate position where herding-induced volatility may provide liquidity opportunities that partially offset adverse selection costs.
Implications for Market Design: This segmentation suggests regulatory interventions should be size-dependent. Circuit breakers or herding monitoring may be most valuable for small-cap stocks where efficiency breakdown is severe, while large-cap stocks may self-correct through arbitrage. Liquidity provision programs targeting small-cap stocks during high herding periods could help maintain price discovery efficiency across the full cross-section. The mechanism analysis (Section 4.4), though limited in explanatory power, is suggestive that volume-based interventions (e.g., circuit breakers, position limits) may warrant more attention than spread-based protections (e.g., minimum tick sizes, market maker obligations).
5.3 Non-Equilibrium Transient Dynamics
The stationarity tests in Table 3 reveal that the Hawkes process governing retail herding has not reached equilibrium, with important implications for interpretation and for the broader market criticality literature.
The apparent paradox—a statistically subcritical branching ratio yet critical-like scaling (Hurst exponent H=0.817H=0.817)—is resolved by recognizing the system is in a transient regime. In systems approaching criticality, critical scaling properties (power-law correlations, long memory) can emerge before the critical point is reached. This is analogous to “critical slowing down” in phase transition theory, where relaxation timescales diverge as the system nears the transition.
The yearly evolution of the branching ratio reveals a clear upward trend:
  • 2020: BR = 0.000 (18 events over 248 days)
  • 2021: BR = 0.420 (16 events over 248 days)
  • 2022: insufficient surge events for reliable estimation
  • 2023: BR = 0.178 (21 events over 245 days)
  • 2024: BR = 0.194 (41 events over 244 days)
Whether this trend continues toward true criticality or stabilizes at some subcritical equilibrium remains an open empirical question.
The transient nature of these dynamics also explains the large discrepancy between the theoretical steady-state intensity ($\bar{\lambda}_{\infty}=\mu/(1-\text{BR})$) and the observed empirical rate noted in Section 4.2. For near-critical processes, the equilibration timescale diverges, and our sample provides insufficient time for the system to reach steady-state.
Economic Interpretation: The non-equilibrium finding suggests that the market’s propensity for retail herding is an evolving state variable, not a stable structural parameter. The post-2020 period coincides with unprecedented shifts in retail market participation: the COVID-19 pandemic, zero-commission mobile trading platforms, the “Donghak ant” retail investor movement, and social media coordination of trading. Our parameter estimates capture this transient period of adjustment, and the upward trend in the branching ratio suggests the market may be undergoing a regime shift toward higher endogenous feedback.
Methodological Implication: All results in this paper should be interpreted as applying to the January 2020 – February 2025 transient regime. Extrapolation to long-run steady-state behavior would require additional assumptions about eventual convergence that our data cannot support. If the upward trend continues, the market could reach true criticality, representing a qualitatively different regime. Monitoring the continued evolution of the branching ratio in out-of-sample periods will be crucial.
Connection to Self-Organized Criticality: The observed approach toward criticality resonates with theories of self-organized criticality (Bak et al.,, 1987), which posit that complex systems naturally evolve toward critical states without external fine-tuning. However, our results remain agnostic on whether this trend represents self-organization or exogenous forcing (e.g., technological or regulatory changes). Distinguishing these mechanisms requires theoretical modeling beyond our empirical scope.
@@ 5.4 Practical Implications
  • Develop real-time monitoring of branching ratios for retail trading activity; high values (approaching 1) signal elevated systemic feedback risk.
  • Interventions should be size-dependent and volume-targeted, as discussed in Section 5.2.
@@ 5.5 Limitations and Future Research
Causality Not Established: While we document strong associations between retail herding intensity and institutional impact deterioration, formal Granger causality tests fail to reject the null (F = 0.020, p = 0.888). The relationships we observe are correlational; the deterioration during herding may reflect common responses to unobserved factors. Future work could explore alternative identification strategies (e.g., instrumental variables, natural experiments) to establish causal mechanisms.
Non-Stationarity and Out-of-Sample Failure: Time-series cross-validation reveals that our deconvolution model does not generalize out-of-sample (Appendix A.6), attributable to the non-stationarity documented in Table 3. All parameter estimates reflect transient dynamics, not equilibrium properties (Section 5.3), and our results should be interpreted as descriptive of the January 2020 – February 2025 period rather than predictive of future behavior.
External Validity: Our sample (January 2020 – February 2025) coincides with extraordinary market conditions—the COVID-19 pandemic, mobile trading platforms, and the “Donghak ant” movement—and focuses on a single market (Korea). While Korea offers excellent investor classification data, the structural parameters we estimate may be specific to this era and geography. Replication on earlier periods (pre-2020) or markets with different institutional structures (e.g., more algorithmic trading in the U.S.) would strengthen generalizability.
Frequency Limitations: We analyze daily frequency data to avoid microstructure noise. However, intraday dynamics may differ. High-frequency herding episodes (e.g., within-day flash crashes) could exhibit even more extreme size-dependent patterns. Future work should extend the regime-conditional deconvolution framework to intraday horizons.
Micro-Foundations and Regime Classification: The precise micro-foundations of size-dependent deterioration remain an open question—whether it arises from differential adverse selection, inventory constraints, or attention constraints across firm sizes could be explored through agent-based models. Additionally, our regime classification treats herding as endogenously determined (Hawkes model), but some episodes may be triggered by exogenous news events. Distinguishing exogenous from endogenous herding triggers would enrich the analysis.
@@ 6 Conclusion
This paper demonstrates that price discovery is not a mechanical constant but a state-dependent phenomenon governed by the dynamic interplay of information and noise.
That the investor-type segmentation documented in Section 4.1 emerges consistently from three independent analytical frameworks strengthens the case that market participant heterogeneity is a first-order feature of price discovery, not a nuisance to be averaged away.
The regime-dependent breakdown during retail herding reveals that efficiency is fragile in precisely the parts of the market least equipped to absorb noise: smaller, less liquid firms. That large-cap stocks maintain and even amplify efficient price discovery during the same episodes underscores that market-wide efficiency metrics can mask severe cross-sectional deterioration.
The relationships we document are correlational, the system operates in a non-equilibrium transient regime that limits out-of-sample generalization, and our results characterize the specific January 2020 – February 2025 period of unprecedented retail market participation. Whether the observed trend toward greater endogenous feedback continues or stabilizes remains an open empirical question.